OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 6(x − 2) double arrow 6x − 12

OpenStudy (anonymous):

what property is it talking about?

OpenStudy (anonymous):

The property is distributive property of multiplication over subtraction. It states that a(b - c) = ab - ac. You see the similarity here?

OpenStudy (anonymous):

6(x - 2) = 6*x - 6*2 = 6x - 12 Hope that helped.

OpenStudy (unklerhaukus):

\[6(x − 2) \Longleftrightarrow 6x − 12 \]

OpenStudy (unklerhaukus):

pratu043 is right it is the distributive property

OpenStudy (anonymous):

its a vector space property 8 which is multiplication over subtraction